### v2

@@ 4.3 Main Results
In Figure 4 and Figure 5, Lemur shows the robust performance in GA and FGA.The consistency and robustness of Lemur are evident in its high FGA and GA across diverse datasets such as HDFS, Hadoop, Spark, etc. This consistency emphasizes its adaptability and robustness to various log parsing requirements. In conclusion, as an unsupervised LLM-related model, Lemur significantly stands out in the log parsing domain. It not only showcases the potential and effectiveness of large language models in unsupervised learning scenarios but also firmly establishes itself as a formidable competitor to both existing supervised and unsupervised models, especially in scenarios where reliance on annotated data is not feasible.
Figure 6 presents a comprehensive analysis of the execution times for multiple algorithms: Lemur, Brain, Drain, IPLOM, Spell, and ULP, across various datasets comprising 2K data points each. This figure provides critical insights into the efficiency and scalability of these algorithms in processing data. Notably, the execution time of Lemur stands out for its relative brevity across the majority of the datasets. This is particularly evident when comparing the performance of the LEMUR algorithm, a component of Lemur, against other baseline methodologies. LEMUR demonstrates a significantly reduced execution time on several key datasets, including Hadoop, BGL, Windows, Android, and Mac, highlighting its efficiency and optimized processing capabilities. Furthermore, we extend this analysis to evaluate the performance scalability of these algorithms across four distinct datasets: BGL, HDFS, Android, and Spark. This extension, as depicted in Figure 7, encompasses varying dataset sizes, thereby providing a more nuanced understanding of each algorithm’s adaptability and performance under different data volume conditions. This comparative analysis is crucial for assessing the practical applicability of these algorithms in real-world scenarios where dataset sizes can vary significantly. The results from this extended analysis are instrumental in drawing conclusions about the overall efficiency and suitability of each algorithm for large-scale data processing in diverse contexts.
5 Ablation
5.1 Effect of Entropy Sampling
Method Average FGA Average PA
@@ 5.2 Effect of Chain-of-Thought Merging
As illustrated in Table 5, which provides a detailed comparison between two versions of Lemur: one implementing the three-hop Chain-of-Thought approach and the other without it. The left columns of the table present the performance metrics of Lemur devoid of COT, while the right columns display the outcomes following the integration of the COT methodology. The data delineates a discernible, albeit modest, improvement in the performance of Lemur when augmented with the COT approach. This enhancement is particularly evident in the FGA and GA metrics across a range of datasets, including ”Hadoop,” ”BGL,” ”HPC,” ”Thunderbird,” ”Linux,” ”Android,” and ”Mac.”
@@ 6 Conclusion
To enhance log analysis in complex software systems, we propose Lemur. This framework replaces traditional rule-based methods with an information entropy-based sampling for efficient log clustering and employs large language models (LLMs) for advanced semantic comprehension in template merging. Lemur has demonstrated superior performance and efficiency in log parsing, validated through extensive tests on large-scale datasets.

### v3

@@ 3.1 Implement Details
3.2 Baselines and Metrics
Baselines
As for baselines, we choose Drain[11], Spell[4], IPLOM[21], ULP[28], Brain[38], LogPPT[18] and LLMParser[15] as our baselines.
F1 score of Grouping Accuracy (FGA)
Grouping Accuracy (GA)
Execution time.
3.3 Main Results
Figure 4: Boxplot of group accuracy on 16 benchmark datasets.
Figure 6: Execution time for various datasets.
Figure 7: Execution time for various dataset sizes.
Method Avg. FGA Avg. GA
@@ 3.4 Effect of Chain-of-Thought Merging
Configuration Avg.FGA Avg.GA
Default (Our Method) 0.999 0.996
KMeans Clustering [30] 0.685 0.698
DBSCAN Clustering [17] 0.731 0.744
Our Method w/o Bucket Grouping 0.355 0.394
Table 6: Additional Ablation Studies on Lemur
In Table 3, Lemur, an unsupervised LLM-related model, exhibits remarkable effectiveness and performance in the field of log parsing. When compared to Brain, the current state-of-the-art in unsupervised models, Lemur demonstrates superior or comparable results across various datasets, underscoring its efficacy in unsupervised learning without the need for annotated data. Moreover, even when juxtaposed with supervised models like LogPPT and LLMParser, Lemur shows equal or better performance in most datasets, based on FGA and GA metrics. This is particularly noteworthy as it highlights the capability of Lemur to match or exceed the performance levels of supervised models, despite the absence of explicit label guidance.
In Figure 5 and Figure 4, Lemur shows the robust performance in GA and FGA. The consistency and robustness of Lemur are evident in its high FGA and GA across diverse datasets such as HDFS, Hadoop, Spark, etc. This consistency emphasizes its adaptability and robustness to various log parsing requirements. In conclusion, as an unsupervised LLM-related model, Lemur significantly stands out in the log parsing domain.
Figure 6 presents a comprehensive analysis of the execution times for multiple algorithms: Lemur, Brain, Drain, IPLOM, Spell, and ULP, across various datasets, comprising 2K data points each. This figure provides critical insights into the efficiency and scalability of these algorithms in processing data. Notably, the execution time of Lemur stands out for its relative brevity across the majority of the datasets. LEMUR demonstrates a significantly reduced execution time on several key datasets, including Hadoop, BGL, Windows, Andriod, and Mac, highlighting its efficiency and optimized processing capabilities. Furthermore, we extend this analysis to evaluate the performance scalability of these algorithms across four distinct datasets: BGL, HDFS, Android, and Spark. This extension, as depicted in Figure 7, encompasses varying dataset sizes, thereby providing a more nuanced understanding of each algorithm’s adaptability and performance under different data volume conditions in real-world scenarios where dataset sizes can vary significantly.
Compared to all unsupervised methods, we have achieved more superior performance in the preliminary stage before employing LLM. Furthermore, compared to other LLM-based methods, which are mostly supervised, such as LogPPT and LLMParser, which are essentially inference after fine-tuning, the training process consumes more GPU resources for local deployment or involves larger data transmission over the network for closed-source ChatGPT. LEMUR achieves better results without fine-tuning the LLM. Whether for local deployment inference or remote API, it only utilizes GPU resources for inference, not for fine-tuning. Thus, it requires fewer GPU resources or less network latency in data transmission. In summary, LEMUR has achieved superior performance and unique advantages for unsupervised log parsing or supervised fine-tuning of LLM.
4 Ablation
4.1 Effect of Entropy Sampling
As illustrated in Table 5, which provides a detailed comparison between two versions of Lemur: one implementing the three-hop Chain-of-Thought approach and the other without it. The left columns of the table present the performance metrics of Lemur devoid of Chain-of-Thought (CoT), while the right columns display the outcomes following the integration of the CoT methodology. The data delineates a useful, albeit modest, improvement in the performance of Lemur when augmented with the CoT approach. This enhancement is particularly evident in the FGA and GA metrics across a range of datasets, including Hadoop, BGL, HPC, Thunderbird, Linux, Android, and Mac.
4.2 Additional Ablation Studies on Lemur
To further validate the robustness of Lemur, we conducted ablation studies focusing on the impact of different clustering algorithms and the effectiveness of the bucket grouping strategy. Specifically, we experimented with replacing our clustering mechanism with KMeans [30] and DBSCAN [17] and evaluated the performance when disabling the bucket grouping based on the length assumption. These modifications aim to probe the sensitivity of Lemur to these components.
The results, as illustrated in Table 6, highlight the critical role of the selected clustering algorithm and bucket grouping strategy in Lemur’s performance. The substantial decrease in Avg.FGA and Avg.GA when employing alternative clustering methods or omitting the bucket grouping underscores their importance in achieving optimal outcomes. This ablation study confirms the necessity of careful component selection and validates the robustness of Lemur’s methodology.
5 Related Work
5.1 Log Parser
In the evolving field of automatic log analysis, crucial for distributed systems and cloud computing, significant progress has been made in log parsing techniques, categorized into frequent pattern mining, clustering, and heuristics rules. Frequent pattern mining is exemplified by SLCT [32] which groups logs based on token frequency, and LogCluster [33] which removes positional constraints in log grouping. Clustering approaches include LogMine [9] with its multi-layered clustering system, LKE[6] using edit distance and position weighing, SHISO [22] improving efficiency through hierarchical clustering, LenMa [29] employing token length vectors, and LPV [37] which uses semantic vectors from word2vec. In the heuristics rules category, IPLOM [21] partitions logs by length and token position, Spell [4] approaches parsing as the longest common sequential problem, Drain [11] groups logs by length and prefixes for template updates, and Prefix-Graph [3] merges prefix trees into graphs for template generation. Recent advancements have introduced deep learning-based algorithms like Nulog [24] Uniparser [20], and LogAP [27] utilizing comparative learning and machine translation for parsing. However, these deep learning methods face challenges in efficiency and high operational costs due to GPU requirements.
5.2 Large Language Model
Language modeling using self-supervised learning and large-scale data, significantly enhances various natural language processing tasks. Specifically, pre-training a Transformer decoder[25, 34] aids in unconditional text generation. Performance improvements in diverse tasks have been linked to the enlargement of Pre-training Language Models (PLMs) by increasing model or data size. This has led to the creation of increasingly larger PLMs, such as GPT-3 with 175 billion parameters and PaLM with 540 billion [2], guided by the scaling laws of large language models [1]. Despite their similar architectures and pre-training tasks, larger PLMs, such as GPT-4 [25], exhibit unique behaviors and emergent abilities, excelling in complex tasks. A prime example is ChatGPT, adapting GPT-series LLMs for engaging dialogues, and showcasing advanced conversational skills. Fine-tuning LLMs on various datasets [36] yields promising results, using human or LLM-created prompts for instruction tuning and refining generations. Chain-of-thought prompting [36], where models explain their reasoning for complex problems, and RLHF [26], a strategy significantly enhance their performance.
@@ 6 Conclusion
To enhance log analysis in complex software systems, we propose Lemur. This framework replaces traditional rule-based methods with an information entropy-based sampling for efficient log clustering and employs large language models (LLMs) for advanced semantic comprehension in template merging. Information entropy streamlines the process of distinguishing between disparate log messages based on their inherent informational content for efficient log clustering. Lemur has demonstrated superior performance and efficiency in log parsing, validated through extensive tests on large-scale datasets.